The measure of an angle is one third the measure of its supplement. What is the measure of the angle?
Answer by rothauserc(4718)

let x be the angles supplement
:
x/3 +x = 180
:
x +3x = 540
:
4x = 540
:
x = 135
:
**********************************************
the measure of the angle is 135/3 = 45 degrees
**********************************************
: